# 2Universität Heidelberg, Interdisziplinäres Zentrum für Wissenschaftliches Rechnen, Heidelberg, Germany

### v3

@@ 2.2 Migration torques
and also accounts for the Lindbland torque $\Gamma_{\mathrm{L}}$ as
@@
The thermal torques are computed as (Masset 2017)
$\Gamma_{\mathrm{th}}=1.61\frac{\gamma-1}{\gamma}\frac{x_{c}}{\lambda}\left(\frac{L}{L_{c}}-1\right)\frac{1}{h}\Gamma_{0},$ (10)
where $x_{c}$ is the corotational radius of the embedded body and $\lambda$ is the typical size of overdense regions surrounding embedded BHs, evaluated as in G24:
@@ 2.4 Nth-generation BHs
For Nth-generation ($Ng$) BHs, which are the remnants of earlier mergers in the AGN disk, we estimate their orbital starting conditions using the outputs of $1g$ simulations. Hence, in the $Ng$ simulations, the initial position $R_{\mathrm{init}}$ of the primary BH is no longer drawn randomly across the disk. Instead, it is sampled directly from the set of $R_{\mathrm{pair}}$ values obtained in the $1g$ simulation, accounting for the natal velocity kick $v_{\mathrm{kick}}$ that is imparted during the merger event. The recoil velocity $v_{\mathrm{kick}}$ is drawn from a precomputed fastcluster catalog (Vaccaro et al. 2024) that already incorporates the dependence on progenitor mass ratio and spin configuration as in Maggiore 2018. The kick direction is randomized assuming isotropic ejection, and contributions from enclosed gas and NSC members are neglected given their small mass relative to the SMBH. If the kick magnitude exceeds the local escape velocity, the remnant is assumed to leave the system. Otherwise, the object is recaptured and is then able to migrate in the disk following the same torque prescriptions as $1g$ BHs. The shifted radius $R_{\mathrm{init}}^{\mathrm{new}}$ is computed by means of simple orbital transfer calculations (Hohmann 1960) as
@@ 3.5 Migration timescales
To assess how well either of these approximations tracks the actual dynamical evolution of BHs in the disk, we compare them to the pair-up times obtained directly from our simulations. These pair-up times, shown via the $40\%$ and $68\%$ contour levels in Figure 7, represent the elapsed time between the moment at which the single primary BH becomes embedded in the disk and the subsequent formation of a BBH. Unlike $t_{\mathrm{migr}}$, they are measured durations rather than derived timescales, and naturally include the effects of stochasticity, mass differences, and local disk structure. Several important features emerge from this comparison. There is substantial scatter in the pair-up times at fixed radius, due in part to the variation in the mass of the migrating BH. There is a sharp upper limit at $t_{\mathrm{migr}}=2$ Myr, corresponding to the assumed AGN disk lifetime adopted in our model. Moreover, some radial regions are systematically disfavored for binary formation, as also noticeable in Figure 6, due to either negative torques in the innermost regions of the disk or to reduced differential migration in the outer disk regions for large $M_{\bullet}$, as discussed in subsection 3.1. Finally, pair-up times in our simulations are typically shorter than the estimated migration timescale in regions where the $\Gamma$ profile has a large slope, and longer where the profile is flatter. This is a consequence of differential migration: pair-ups are set by relative drift rather than absolute torque. Where the migration speed profile is steep, nearby objects converge quickly and pair up fast. Vice versa, where the profile is flat, velocity differences are only mass-dependent and consequently pair-up times lengthen.
@@ 4 Caveats
Our treatment of Type II migration is deliberately conservative, as gap-opening objects are assumed to effectively prevent any inflow across the gap. More realistically, the transition to Type II is not a hard switch: the gap would not open fully and gas inflow would not be completely suppressed (Malik et al. 2015). As a result, the Type II regime would more closely resemble a modified Type I migration with reduced local density (as shown in Kanagawa et al. 2018 for $\alpha\leq 0.01$). We explore this alternative model for Type II migration in Appendix B.
Moreover, our model does not include resonant interactions, such as mean-motion resonances, which could alter or halt migration. We also do not consider dynamical formation via three-body scatterings, which is another viable mechanism of BBH formation in AGN environments (Tagawa et al. 2020b). Finally, while our pair-up criterion identifies candidate binaries based on proximity, it does not explicitly model binary binding or survivability under the influence of gas torques and disk turbulence.
@@ Appendix B Type II migration
where the contrast parameter $K$ is defined as in eq. 13. For low-mass migrators ($K\lesssim 1$, $\Sigma_{\mathrm{g}}^{\prime}\gtrsim 0.96\,\Sigma_{\mathrm{g}}$), the gap is shallow and standard Type I migration (eq. 1) remains valid. As the migrator mass increases, the contrast parameter $K$ grows quadratically. The full transition to a Type II regime happens for $K\gtrsim 20$ (K18), as the disk surface density at the orbit is significantly reduced to $\Sigma_{\mathrm{g}}^{\prime}\lesssim 0.55\,\Sigma_{\mathrm{g}}$. In this regime, both the normalization of the torque $\Gamma_{0}$ (eq. 12), and the shape of the torque profiles (via the parameter $a$ in eq. 4) are significantly modified.
We stress that the K18 prescription has only been calibrated for low viscosities ($\alpha\leq 0.01$) and moderate aspect ratios ($0.03\lesssim h\lesssim 0.07$). Its applicability at $\alpha>0.01$, typical of some AGN disk models, remains uncertain. For this reason, in the main body of this paper we adopt the classical Type II migration prescription, while here we outline the implications of the Kanagawa et al. (2018) model.
@@
The fraction of pair-ups occurring within migration traps, $f_{\mathrm{trap}}$, retains a similar behavior than in Figure 3 for the G24 case, remaining large (>0.6>0.6) for most of the parameter space, and decreasing for SMBH masses above $10^{8.4}\,\mathrm{M}_{\odot}$. Contrarily, in the B16 case traffic-jam pair-ups are more relevant, especially at low viscosity for low SMBH mass. Therefore, $f_{\mathrm{trap}}<0.4$ with the exception of a certain $\alpha$-dependent window: $\log M_{\bullet}/\,\mathrm{M}_{\odot}\in\left[6.6,8.5\right]$ for $\alpha=0.01$, $\log M_{\bullet}/\,\mathrm{M}_{\odot}\in\left[6.0,7.5\right]$ for $\alpha=0.1$, or $\log M_{\bullet}/\,\mathrm{M}_{\odot}\in\left[5.5,7.2\right]$ for $\alpha=0.4$. It retains a similar trend in both migrator mass $m$ and initial radius $R_{\mathrm{init}}$ than discussed in subsection 3.4, with the exception

### v4

@@ 2.2 Migration torques
and also accounts for the Lindblad torque $\Gamma_{\mathrm{L}}$ as
@@
The thermal torques are computed as (Masset 2017; Velasco Romero & Masset 2020)
$\Gamma_{\mathrm{th}}=1.61\frac{\gamma-1}{\gamma}\frac{x_{c}}{\lambda}\left(\frac{L}{L_{c}}\,\frac{4m_{c}}{m+4m_{c}}-\frac{2m_{c}}{m+2m_{c}}\right)\frac{\Gamma_{0}}{h}\left(1-e^{-\lambda\tau_{V}/h}\right),$ (10)
where $m_{c}=c_{s}\chi/G$ is a critical mass, $\tau_{V}$ is the vertical optical depth, $x_{c}$ is the corotational radius of the embedded body and $\lambda$ is the typical size of overdense regions surrounding embedded BHs, evaluated as in G24:
@@ 2.4 Nth-generation BHs
For Nth-generation ($Ng$) BHs, which are the remnants of earlier mergers in the AGN disk, we estimate their orbital starting conditions using the outputs of $1g$ simulations. Hence, in the $Ng$ simulations, the initial position $R_{\mathrm{init}}$ of the primary BH is no longer drawn randomly across the disk. Instead, it is sampled directly from the set of $R_{\mathrm{pair}}$ values obtained in the $1g$ simulation, accounting for the natal velocity kick $v_{\mathrm{kick}}$ that is imparted during the merger event. The recoil velocity $v_{\mathrm{kick}}$ is drawn from a precomputed fastcluster catalog (Vaccaro et al. 2024) that already incorporates the dependence on progenitor mass ratio and spin configuration as in Maggiore (2018). The kick direction is randomized assuming isotropic ejection, and contributions from enclosed gas and NSC members are neglected given their small mass relative to the SMBH. If the kick magnitude exceeds the local escape velocity, the remnant is assumed to leave the system. Otherwise, the object is recaptured and is then able to migrate in the disk following the same torque prescriptions as $1g$ BHs. The shifted radius $R_{\mathrm{init}}^{\mathrm{new}}$ is computed by means of simple orbital transfer calculations (Hohmann 1960) as
@@ 3.5 Migration timescales
To assess how well either of these approximations tracks the actual dynamical evolution of BHs in the disk, we compare them to the pair-up times, $t_{\mathrm{pair}}$, obtained directly from our simulations. These pair-up times, shown via the $40\%$ and $68\%$ contour levels in Figure 7, represent the elapsed time between the moment at which the single primary BH becomes embedded in the disk and the subsequent formation of a BBH. Unlike $t_{\mathrm{migr}}$, they are measured durations rather than derived timescales, and naturally include the effects of stochasticity, mass differences, and local disk structure. Several important features emerge from this comparison. There is substantial scatter in the pair-up times at fixed radius, due in part to the variation in the mass of the migrating BH. There is a sharp upper limit at $t_{\mathrm{migr}}=2$ Myr, corresponding to the assumed AGN disk lifetime adopted in our model. Moreover, some radial regions are systematically disfavored for binary formation, as also noticeable in Figure 6, due to either negative torques in the innermost regions of the disk or to reduced differential migration in the outer disk regions for large $M_{\bullet}$, as discussed in subsection 3.1. Finally, pair-up times in our simulations are typically shorter than the estimated migration timescale in regions where the $\Gamma$ profile has a large slope, and longer where the profile is flatter. This is a consequence of differential migration: pair-ups are set by relative drift rather than absolute torque. Where the migration speed profile is steep, nearby objects converge quickly and pair up fast. Vice versa, where the profile is flat, velocity differences are only mass-dependent and consequently pair-up times lengthen.
@@ 4 Caveats
Our treatment of Type II migration is deliberately conservative, as gap-opening objects are assumed to effectively prevent any inflow across the gap. More realistically, the transition to Type II is not a hard switch: the gap would not open fully and gas inflow would not be completely suppressed (Malik et al. 2015). As a result, the Type II regime would more closely resemble a modified Type I migration with reduced local density (as shown in Kanagawa et al. 2018; Fung et al. 2014). We explore this alternative model for Type II migration in Appendix B.
Moreover, our model does not include resonant interactions, such as mean-motion resonances, which could alter or halt migration (Epstein-Martin et al. 2025; Moncrieff et al. 2025). We also do not consider dynamical formation via three-body scatterings, which is another viable mechanism of BBH formation in AGN environments (Tagawa et al. 2020b). Finally, while our pair-up criterion identifies candidate binaries based on proximity, it does not explicitly model binary binding or survivability under the influence of gas torques and disk turbulence.
@@ Appendix B Type II migration
Figure 12: Pair-up times $t_{\mathrm{pair}}$ as a function of SMBH mass in SG AGN disks, for different combinations of the viscosity parameter $\alpha$ (left: 0.01; right: 0.1) and torque prescriptions (top: G24; bottom: B16). Each panel shows a median line and contours enclosing $68\%$ and $90\%$ of the pair-up time distribution, based on our simulations. Results are shown for two different Type II migration prescriptions, indicated by color and line style. In blue, dashed lines: full gap opening, as in the main text (eq. 15). In red, continuous lines: partial gap opening, as in this appendix (K18).
@@
where the contrast parameter $K$ is defined as in eq. 13. For low-mass migrators ($K\lesssim 1$, $\Sigma_{\mathrm{g}}^{\prime}\gtrsim 0.96\,\Sigma_{\mathrm{g}}$), the gap is shallow and standard Type I migration (eq. 1) remains valid. As the migrator mass increases, the contrast parameter $K$ grows quadratically. The full transition to a Type II regime happens for $K\gtrsim 20$ (K18), as the disk surface density at the orbit is significantly reduced to $\Sigma_{\mathrm{g}}^{\prime}\lesssim 0.55\,\Sigma_{\mathrm{g}}$.
We calculate corotational gas torques in eq.s 3, 4, 5, 6, 7, and 12 using the reduced local density $\Sigma_{\mathrm{g}}^{\prime}$. In this regime, both the normalization of the torque $\Gamma_{0}$ (eq. 12), and the shape of the torque profiles (via the parameter $a$ in eq. 4) are significantly modified. We leave the Lindblad torque ($\Gamma_{\mathrm{L}}$, eq. 8) unperturbed.
The effect of gap opening of Lindblad torques is quite complex, as it depends on the exact profile of the gap (e.g. Kley & Nelson 2012; Petrovich & Rafikov 2012). We performed tests in which we rescaled them with the density at the bottom of the gap, $\Sigma^{\prime}_{g}$, as we did with corotational torques, with minimal changes to our results. A more comprehensive simulation would include modeling the gap’s density profile and computing Lindblad torques self-consistently. However, in the high-viscosity regime in which we are operating, we do not expect any non-linear behavior that would strongly affect our results (Duffell 2015).
@@
The fraction of pair-ups occurring within migration traps, $f_{\mathrm{trap}}$, retains a similar behavior than in Figure 3 for the G24 case, remaining large (>0.6>0.6) for most of the parameter space, and decreasing for SMBH masses above $10^{8.4}\,\mathrm{M}_{\odot}$. Contrarily, in the B16 case traffic-jam pair-ups are more relevant, especially at low viscosity for low SMBH mass. Therefore, $f_{\mathrm{trap}}<0.4$ with the exception of a certain $\alpha$-dependent window: $\log M_{\bullet}/\,\mathrm{M}_{\odot}\in\left[6.6,8.5\right]$ for $\alpha=0.01$, $\log M_{\bullet}/\,\mathrm{M}_{\odot}\in\left[6.0,7.5\right]$ for $\alpha=0.1$, or $\log M_{\bullet}/\,\mathrm{M}_{\odot}\in\left[5.5,7.2\right]$ for $\alpha=0.4$. It retains a similar trend in both migrator mass $m$ and initial radius $R_{\mathrm{init}}$ than discussed in subsection 3.4.
We compute pair-up times, $t_{\mathrm{pair}}$, and evaluate how the K18 framework affects them. These pair-up times, shown via the $68\%$ and $90\%$ contour levels in Figure 12, represent the elapsed time between the moment at which the single primary BH becomes embedded in the disk and the subsequent formation of a BBH. At low SMBH mass, $M_{\bullet}\lesssim 3\times 10^{6}\,\mathrm{M}_{\odot}$, the normalized contrast parameter can reach larger values, $K_{\mathrm{norm}}\gtrsim 1$, thus significantly reducing the local gas density $\Sigma^{\prime}$ and slowing down the pair-up process, therefore increasing $t_{\mathrm{pair}}$ by almost two orders of magnitude for the lowest SMBH mass explored, $M_{\bullet}=10^{5}\,\mathrm{M}_{\odot}$. In contrast, for $M_{\bullet}\gtrsim 3\times 10^{6}\,\mathrm{M}_{\odot}$, pair-up times are unaffected by the adopted Type II prescription.